4 men or 6 boys can dig a well in 14 days . how long would 2 men and 4 boys take to dig the same well?
Answer by ankor@dixie-net.com(22740)

4 men or 6 boys can dig a well in 14 days.
Find the number of man-days and boy-days required to complete the job
4 * 14 = 56 man-days
6 * 14 = 84 boy-days
:
Since they both worked 14 days to complete the job, we can say:
4m = 6b
2m = 3b
m = 1.5b
:
how long would 2 men and 4 boys take to dig the same well?
Use boy-days, replace men with 1.5
let d = no. of days required with this scenario
2(1.5)d + 4d = 84
3d + 4d = 84
7d = 84
d = 84/7
d = 12 days for 2 men and 4 boys to do the job